# DCT-Based Decorrelated Attention for Vision Transformers

### v2

@@ Abstract
1 Introduction
In the last decade, convolutional neural networks (CNNs) have dominated image-processing tasks. However, inspired by the success of transformer architectures in natural language processing, Vision Transformers (ViTs) in recent years have redefined the norms of visual data interpretation. Self-attention mechanism, a key component of transformers, is pivotal in reshaping how models perceive and analyze visual information [49]. Unlike CNNs, ViTs view an image as a sequence of informative patches, adopting a more holistic approach. By treating image patches like “visual words”, ViTs can leverage powerful transformer architectures, originally designed for natural language processing, to learn global context and long-range dependencies. This novel approach has led to remarkable performance gains on a wide range of computer vision tasks, demonstrating the potential of ViTs to revolutionize the field.
@@ 1 Introduction
Efficiency is an ongoing concern in Transformers. While conventional initialization strategies are often suitable for generic applications, new algorithms are increasingly tailored to the specific characteristics and needs of Transformers. These advancements in initialization techniques contribute to the ongoing improvements in the performance and efficiency of Transformer-based models. Orthogonal initialization [40, 21], sparse initialization [31], variance scaling, pre-trained initialization [51, 4], layer-wise, domain-specific and attention-specific initialization [48] are some of the new trends in initialization approaches. In this study, we propose a new family of algorithms for initialization strategy in the intersection of attention and domain-specific approaches: discrete cosine attention.
@@
2 Related works
@@
Vision Transformers Transformer architectures have merged as a dominant standard for natural language processing (NLP) tasks [49, 53]. Vision Transformer (ViT) [14] built a bridge between language and vision. It shows that pure transformers can also be applied directly to sequences of image patches for classification tasks. In other words, it partitions images into patches and treats them as word sequences. Then, a shifted windowing scheme was proposed in Swin Transformer [29] to obtain greater efficiency. This limits self-attention computation to non-overlapping local windows while allowing for cross-window connection. Later, to overcome the limitations of the heavyweights in ViTs, MobileViT [33] combined the strengths of MobileNets [39] and ViTs. It is a lightweight and low-latency network for mobile vision tasks. Moreover, gSwin [17] combined MLP-based architecture with Swin Transformer and attains superior accuracy with a smaller model size. Vision transformers are now supreme in various computer vision tasks such as super resolution [27, 57], compressive sensing [43, 54], image segmentation [59, 23, 47], and human mesh reconstruction [62].
@@
Frequency-based Neural Networks Orthogonal transforms such as Fourier transform [10, 5], Hadamard transform [34, 36], and discrete cosine Stockwell transform [64] have been used to improve the performance of the neural networks. In the orthogonal transform family, DCT is a mathematical technique commonly used in signal processing and image and video compression [35]. Compared to the Fourier transform, which is the standard frequency representation, DCT does not generate complex-valued numbers. Therefore, it is computationally more efficient. DCT-Former [41], for instance, applied DCT compression on the input sequence along the length channel before feeding into the self-attention layers to reduce the computational overhead in the NLP problems. Furthermore, inspired by JPEG compression, DCFormer [26] highlighted the visually significant signals by leveraging the input information compression on its frequency domain representation. These two DCT-based works and our DCT-based compression all aim to reduce the computational overhead. Therefore, we mainly compare our DCT-based compression method with them. DCT-Former [41] is designed for the NLP transformers and takes DCT along the width and height of the input tensor patches, while our method is designed for the vision transformers and it compresses the tensor along the channel axis. DCFormer [26] averages DCT coefficients via average pooling, which does not take advantage of the property that DCT components in low-frequency contain more important information. On the contrary, our DCT-based compression method retains the low-frequency components very accurately and discards the unnecessary extremely high-frequency components corresponding to noise. Under the same input image size, compared to the vanilla transformer models, DCFormer reduces the FLOPs and retains the number of parameters while sacrificing accuracy. Our DCT-based compression method, on the other hand, saves both trainable parameters and FLOPs and obtains on-par or better performance in addition to other benefits that our DCT-based compression method has. Furthermore, we introduce a DCT-based initialization and we have not found any related initialization in the literature.
Initialization Methods for Self-Attention Mimetic initialization [48] is a compute-free method to initialize the weights of self-attention layers. It simulated the patterns present in the pre-trained weights to achieve performance improvement of the transformers.
@@
Efficient Multi-head Self-Attention To reduce the memory and computational costs of dot-product attention, an efficient mechanism [42] in linear complexity was proposed. Besides, Hydra Attention [3] increased the number of heads to address the issue of the high computational complexity of the self-attention mechanism on large-sized images. Further, MetaFormer [55] verified that the overall architecture of Transformers, rather than the specific token mixer module, plays a more crucial role in the performance.
3 Methodology
3.1 Background: Multi-Head Self-Attention
@@
The overall structure of the vanilla MSA is drawn in Figure 1a. To handle a 2D image tensor $\mathbf{I}\in\mathbb{R}^{H\times\ W\times C}$, where $H\times W$ are the resolution of the image and $C$ is the number of channels, we first split it into patches $\mathbf{X}\in\mathbb{R}^{N\times M^{2}\times C}$, where $M\times M$ is the resolution of each image patch and $N=\frac{WH}{M^{2}}$ is the resulting number of patches. After that, three linear layers ($\mathcal{L}_{Q}$, $\mathcal{L}_{K}$, $\mathcal{L}_{V}$) are applied on $\mathbf{X}$ to compute the queries, keys, and values $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{N\times M^{2}\times C}$:
@@
3.2 DCT-based Initialization for Attention
@@
$\{\mathbf{X}\mathcal{D}^{T}\}[k]=\begin{cases}\frac{\sqrt{2}}{C}\sum_{n=0}^{C-1}\mathbf{X}[n],&k=0,\\ \frac{2}{C}\sum_{n=0}^{C-1}\mathbf{X}[n]\cos\frac{(2n+1)k\pi}{2nC},&k=1,\ldots,C-1.\end{cases}$ (4)
@@
3.3 DCT-Based Compressed Attention
@@
$\{\mathbf{X}\mathcal{D}^{T}\}[k]=\begin{cases}\frac{\sqrt{2}}{C}\sum_{n=0}^{C-1}\mathbf{X}[n],&k=0,\\ \frac{2}{C}\text{Re}\left(e^{-\frac{ik\pi}{2C}}\sum_{n=0}^{2C-1}\mathbf{X}[n]W^{kn}\right),&k=1,\ldots,C-1,\end{cases}$ (5)
@@
We observed that the attention layer matrices are correlated with each other so that we can compress them using DCT. Figure 1b shows the method of DCT-based compression for the attention layer. First, the input patches $\mathbf{X}\in\mathbb{R}^{N\times M^{2}\times C}$ are encoded by DCT along the channel of $C$ with the high-frequency components truncated. If we keep the first $\tau C$ coefficients ($\tau<1$), the essential information of $\mathbf{X}$ is still preserved, while the input and output dimensions of the linear layers ($\mathcal{L}_{Q},\mathcal{L}_{K},$ and $\mathcal{L}_{V}$) are reduced from $\mathbb{R}^{N\times M^{2}\times C}$ to $\mathbb{R}^{N\times M^{2}\times\tau C}$. In this way, the computational cost from $\mathcal{L}_{Q},\mathcal{L}_{K}$, and $\mathcal{L}_{V}$ is reduced significantly. In our experiments, $\tau$ is chosen from 25%, 50%, and 75%. After obtaining the attention $\mathbf{A}\in\mathbb{R}^{N\times M^{2}\times\tau C}$ from the queries $\mathbf{Q}$, keys $\mathbf{K}$, and values $\mathbf{V}$, we apply zero-padding along the last channel to make the dimension back to $\mathbb{R}^{N\times M^{2}\times C}$. Before being fed into the linear layer $\mathcal{L}_{O}$, these padded patches are decoded by the inverse DCT (IDCT):
$\{\mathbf{A}(\mathcal{D}^{-1})^{T}\}[n]=\frac{\sqrt{2}}{2}\mathbf{A}[0]+\sum_{k=1}^{C-1}\mathbf{A}[k]\cos\frac{(2n+1)k\pi}{2C},n=0,1,\ldots,C-1.$ (6)
@@
DCT and IDCT can be implemented in a fast manner with the complexity of $O(C\log_{2}C)$ using butterfly algorithm [9]. However, such an implementation is not officially supported in PyTorch currently. Therefore, in this work, we implement the DCT and IDCT via matrix multiplication between the input tensor and the truncated DCT and IDCT matrices $\bar{\mathcal{D}}\in\mathbb{R}^{\tau C\times C}$ and $\bar{\mathcal{D}}^{-1}\in\mathbb{R}^{C\times\tau C}$. These transform matrices are obtained by truncating the DCT and IDCT matrices along rows or columns to keep $\tau$ coefficients or zero padding, respectively. Then, DCT with truncation or IDCT with zero padding is implemented as one matrix multiplication to reduce memory pressure. Furthermore, as no non-linearity exists between IDCT and the final linear layer $\mathcal{L}_{O}$, these two steps can be implemented as a single linear layer whose weight matrix is $\mathbf{W}_{O}\bar{\mathcal{D}}^{-1}$ and the bias is still $\mathbf{b}_{O}$. However, such a combination is not applied on DCT and $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, or $\mathbf{W}_{V}$. This is because it may not always reduce the computational cost for them: To compute $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$, with such a combination, $3kNM^{2}C^{3}$ multiplications are needed. In contrast, if they are computed separately, there are $\tau NM^{2}C^{3}$ multiplications from DCT and $\tau^{2}NM^{2}C^{3}$ multiplications from each linear layer. As a result, there are a total of $(\tau+3\tau^{2})NM^{2}C^{3}$ multiplications. Hence, combing DCT with $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$ and $\mathbf{W}_{V}$ brings computational cost saving only when $\tau>\frac{2}{3}$. To maintain the consistency of the framework, we do not combine DCT with $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, or $\mathbf{W}_{V}$.
@@
Computational cost comparison of the vanilla versus the proposed DCT-compressed attentions is presented in Table 1. We evaluate both naive and simplified implementations of the DCT-compressed attention with the vanilla attention. The simplification is implemented by fusing IDCT into the linear layer $\mathcal{L}_{O}$. Computational overhead from the Softmax function is omitted because the Softmax function is based on exponential terms. Despite this, the Softmax function in the DCT-compressed attention requires less computational cost than in the vanilla attention as the number of entries in $\left(\frac{\mathbf{QK}^{T}}{\sqrt{d}}+\mathbf{B}\right)$ in the DCT-compressed attention is $\tau^{2}$ times ($\tau<1$) as in the vanilla attention.
@@
Table 1: Multiplications in MSA on $\mathbf{X}\in\mathbb{R}^{N\times M^{2}\times C}$. “--” means this step is not required. $\tau<1$.
@@
Total $4NM^{2}C^{3}$ $(1+2\tau+3\tau^{2})NM^{2}C^{3}$ $(2\tau+3\tau^{2})NM^{2}C^{3}$
$+2NM^{4}C^{2}$ $+2\tau^{2}NM^{4}C^{2}$ $+2\tau^{2}NM^{4}C^{2}$
3.4 Frequency Domain Motivation of the DCT-based Initialization

### v3

# Discrete Cosine Transform Based Decorrelated Attention for Vision Transformers

@@ Abstract
Index Terms:
Discrete cosine transform, attention, vision transforms.
I Introduction
In the last decade, convolutional neural networks (CNNs) have dominated image-processing tasks. However, inspired by the success of transformer architectures in natural language processing, Vision Transformers (ViTs) have redefined the norms of visual data interpretation in recent years. Self-attention mechanism, a key component of transformers, is pivotal in reshaping how models perceive and analyze visual information [1]. Unlike CNNs, ViTs view an image as a sequence of informative patches, adopting a more holistic approach. By treating image patches like “visual words”, ViTs can leverage powerful transformer architectures, originally designed for natural language processing, to learn global context and long-range dependencies. This novel approach has led to remarkable performance gains on a wide range of computer vision tasks, demonstrating the potential of ViTs to revolutionize the field.
@@ I Introduction
Efficiency is an ongoing concern in Transformers. While conventional initialization strategies are often suitable for generic applications, new algorithms are increasingly tailored to the specific characteristics and needs of Transformers. These advancements in initialization techniques contribute to the ongoing improvements in the performance and efficiency of Transformer-based models. Orthogonal initialization [7, 8], sparse initialization [9], variance scaling [10], pre-trained initialization [11, 12, 13], layer-wise, domain-specific and attention-specific initialization [14] are some of the new trends in initialization approaches. In this study, we propose a new family of algorithms for initialization strategy in the intersection of attention and domain-specific approaches: discrete cosine attention.
@@
II Related works
II-A Vision Transformers
@@
Transformer architectures have merged as a dominant standard for natural language processing (NLP) tasks [1, 15]. Vision Transformer (ViT) [16] built a bridge between language and vision. It shows that pure transformers can also be applied directly to sequences of image patches for classification tasks. In other words, it partitions images into patches and treats them as word sequences. Then, a shifted windowing scheme was proposed in Swin Transformer [17] to obtain greater efficiency. This limits self-attention computation to non-overlapping local windows while allowing for cross-window connection. Later, to overcome the limitations of the heavyweights in ViTs, MobileViT [18] combined the strengths of MobileNets [19] and ViTs. It is a lightweight and low-latency network for mobile vision tasks. Moreover, gSwin [20] combined MLP-based architecture with Swin Transformer and attains superior accuracy with a smaller model size. Vision transformers are now supreme in various computer vision tasks such as super resolution [21, 22], compressive sensing [23, 24], image segmentation [25, 26, 27], and human mesh reconstruction [28].
II-B Frequency-based Neural Networks
@@
Orthogonal transforms such as Fourier transform [29, 30], Hadamard transform [31, 32], and discrete cosine Stockwell transform [33] have been used to improve the performance of the neural networks. In the orthogonal transform family, DCT is a mathematical technique commonly used in signal processing and image and video compression [34]. Compared to the Fourier transform, which is the standard frequency representation, DCT does not generate complex-valued numbers. Therefore, it is computationally more efficient. DCT-Former [35], for instance, applied DCT compression on the input sequence along the length channel before feeding into the self-attention layers to reduce the computational overhead in the NLP problems. Furthermore, inspired by JPEG compression, DCFormer [36] highlighted the visually significant signals by leveraging the input information compression on its frequency domain representation. These two DCT-based works and our DCT-based compression all aim to reduce the computational overhead. Therefore, we mainly compare our DCT-based compression method with them. DCT-Former [35] is designed for the NLP transformers and takes DCT along the width and height of the input tensor patches, while our method is designed for the vision transformers and it compresses the tensor along the channel axis. DCFormer [36] averages DCT coefficients via average pooling, which does not take advantage of the property that DCT components in low-frequency contain more important information. On the contrary, our DCT-based compression method retains the low-frequency components very accurately and discards the unnecessary extremely high-frequency components corresponding to noise. Under the same input image size, compared to the vanilla transformer models, DCFormer reduces the FLOPs and retains the number of parameters while sacrificing accuracy. Our DCT-based compression method, on the other hand, saves both trainable parameters and FLOPs and obtains on-par or better performance in addition to other benefits that our DCT-based compression method has. Furthermore, we introduce a DCT-based initialization and we have not found any related initialization in the literature.
II-C Initialization Methods for Self-Attention
DS-Init [10] improves convergence in deep Transformers for NLP by reducing parameter variance during initialization and minimizing output variance of residual connections, facilitating better gradient flow. It pairs with a merged attention sublayer (MAtt) that combines average-based self-attention with encoder-decoder attention to reduce the computational cost. Additionally, Mimetic initialization [14] is a compute-free method that initializes the weights of self-attention layers by simulating the patterns found in pre-trained weights, resulting in improved Transformer performance. Moreover, weight selection [13] enables the initialization of smaller models by selecting a subset of weights from a larger pre-trained model, facilitating efficient knowledge transfer and faster training. This method enhances small model performance, reduces training time, and integrates with knowledge distillation, making it particularly valuable for resource-constrained environments. Furthermore, structured initialization [37] tackles the challenge of training Vision Transformers (ViTs) on small-scale datasets by introducing a structured initialization strategy inspired by the architectural biases of Convolutional Neural Networks (CNNs). While these methods focus on mimicking pre-trained patterns, selecting weights from larger models, or using CNN-inspired biases, our DCT-based initialization utilizes well-established DCT matrices to initialize attention weights.
II-D Efficient Multi-head Self-Attention
@@
To reduce the memory and computational costs of dot-product attention, an efficient mechanism [38] in linear complexity was proposed. Besides, Hydra Attention [39] increased the number of heads to address the issue of the high computational complexity of the self-attention mechanism on large-sized images. Further, MetaFormer [40] verified that the overall architecture of Transformers, rather than the specific token mixer module, plays a more crucial role in the performance. Existing methods reduce self-attention complexity through linear attention, more heads, or architectural optimizations but ignore frequency-domain transformations. We propose DCT-based methods to enhance efficiency and expressiveness in ViTs.
III Methodology
In this section, we first review the multi-head self-attention (MSA) [17]. Then, we introduce our two proposed strategies to utilize the discrete cosine transform (DCT) in the attention function of the Swin Transformers. Finally, we discuss the motivation of our methods.
III-A Background: Multi-Head Self-Attention
@@
The overall structure of the vanilla MSA is drawn in Fig. 1a. To handle a 2D image tensor $\mathbf{I}\in\mathbb{R}^{H\times\ W\times C}$, where $H\times W$ are the resolution of the image and $C$ is the number of channels, we first split it into patches $\mathbf{X}\in\mathbb{R}^{N\times M^{2}\times C}$, where $M\times M$ is the resolution of each image patch and $N=\frac{WH}{M^{2}}$ is the resulting number of patches. After that, three linear layers ($\mathcal{L}_{Q}$, $\mathcal{L}_{K}$, $\mathcal{L}_{V}$) are applied on $\mathbf{X}$ to compute the queries, keys, and values $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{N\times M^{2}\times C}$:
@@
III-B DCT-based Initialization for Attention
@@
$\begin{split}&\{\mathbf{X}\mathcal{D}^{T}\}[k]\\ &=\begin{cases}\frac{\sqrt{2}}{C}\sum_{n=0}^{C-1}\mathbf{X}[n],&k=0,\\ \frac{2}{C}\sum_{n=0}^{C-1}\mathbf{X}[n]\cos\frac{(2n+1)k\pi}{2nC},&k=1,\ldots,C-1.\end{cases}\end{split}$ (4)
@@
III-C DCT-Based Compressed Attention
@@
$\begin{split}&\{\mathbf{X}\mathcal{D}^{T}\}[k]\\ &=\begin{cases}\frac{\sqrt{2}}{C}\sum_{n=0}^{C-1}\mathbf{X}[n],&k=0,\\ \frac{2}{C}\text{Re}\left(e^{-\frac{ik\pi}{2C}}\sum_{n=0}^{2C-1}\mathbf{X}[n]W^{kn}\right),&k=1,\ldots,C-1,\end{cases}\end{split}$ (5)
@@
We observed that the attention layer matrices are correlated with each other so that we can compress them using DCT. Fig. 1b shows the method of DCT-based compression for the attention layer. First, the input patches $\mathbf{X}\in\mathbb{R}^{N\times M^{2}\times C}$ are encoded by DCT along the channel of $C$ with the high-frequency components truncated. If we keep the first $\tau C$ coefficients ($\tau<1$), the essential information of $\mathbf{X}$ is still preserved, while the input and output dimensions of the linear layers ($\mathcal{L}_{Q},\mathcal{L}_{K},$ and $\mathcal{L}_{V}$) are reduced from $\mathbb{R}^{N\times M^{2}\times C}$ to $\mathbb{R}^{N\times M^{2}\times\tau C}$. In this way, the computational cost from $\mathcal{L}_{Q},\mathcal{L}_{K}$, and $\mathcal{L}_{V}$ is reduced significantly. In our experiments, $\tau$ is chosen from 25%, 50%, and 75%. After obtaining the attention $\mathbf{A}\in\mathbb{R}^{N\times M^{2}\times\tau C}$ from the queries $\mathbf{Q}$, keys $\mathbf{K}$, and values $\mathbf{V}$, we apply zero-padding along the last channel to make the dimension back to $\mathbb{R}^{N\times M^{2}\times C}$. Before being fed into the linear layer $\mathcal{L}_{O}$, these padded patches are decoded by the inverse DCT (IDCT):
$\begin{split}\{\mathbf{A}(\mathcal{D}^{-1})^{T}\}[n]=\frac{\sqrt{2}}{2}\mathbf{A}[0]+\sum_{k=1}^{C-1}\mathbf{A}[k]\cos\frac{(2n+1)k\pi}{2C}&,\\ n=0,1,\ldots,C-1&.\end{split}$ (6)
@@
Because $\mathcal{D}$ is orthogonal, i.e. $\mathcal{DD}^{T}=\mathbf{I}_{C\times C}$, the IDCT matrix can be obtained directly from the transpose of the DCT matrix. DCT and IDCT can be implemented in a fast manner with the complexity of $O(C\log_{2}C)$ using butterfly algorithm [42]. However, such an implementation is not officially supported in PyTorch currently. Therefore, in this work, we implement the DCT and IDCT via matrix multiplication between the input tensor and the truncated DCT and IDCT matrices $\bar{\mathcal{D}}\in\mathbb{R}^{\tau C\times C}$ and $\bar{\mathcal{D}}^{-1}\in\mathbb{R}^{C\times\tau C}$. These transform matrices are obtained by truncating the DCT and IDCT matrices along rows or columns to keep $\tau$ coefficients or zero padding, respectively. Then, DCT with truncation or IDCT with zero padding is implemented as one matrix multiplication to reduce memory pressure. Furthermore, as no non-linearity exists between IDCT and the final linear layer $\mathcal{L}_{O}$, these two steps can be implemented as a single linear layer whose weight matrix is $\mathbf{W}_{O}\bar{\mathcal{D}}^{-1}$ and the bias is still $\mathbf{b}_{O}$. However, such a combination is not applied on DCT and $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, or $\mathbf{W}_{V}$. This is because it may not always reduce the computational cost for them: To compute $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$, with such a combination, $3kNM^{2}C^{3}$ multiplications are needed. In contrast, if they are computed separately, there are $\tau NM^{2}C^{3}$ multiplications from DCT and $\tau^{2}NM^{2}C^{3}$ multiplications from each linear layer. As a result, there are a total of $(\tau+3\tau^{2})NM^{2}C^{3}$ multiplications. Hence, combing DCT with $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$ and $\mathbf{W}_{V}$ brings computational cost saving only when $\tau>\frac{2}{3}$. To maintain the consistency of the framework, we do not combine DCT with $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, or $\mathbf{W}_{V}$.
@@
Computational cost comparison of the vanilla versus the proposed DCT-compressed attentions is presented in Table I. We evaluate both naive and simplified implementations of the DCT-compressed attention with the vanilla attention. The simplification is implemented by fusing IDCT into the linear layer $\mathcal{L}_{O}$. Computational overhead from the Softmax function is omitted because the Softmax function is based on exponential terms. Despite this, the Softmax function in the DCT-compressed attention requires less computational cost than in the vanilla attention as the number of entries in $\left(\frac{\mathbf{QK}^{T}}{\sqrt{d}}+\mathbf{B}\right)$ in the DCT-compressed attention is $\tau^{2}$ times ($\tau<1$) as in the vanilla attention.
@@
TABLE I: Multiplications in MSA on $\mathbf{X}\in\mathbb{R}^{N\times M^{2}\times C}$. “--” means this step is not required. $\tau<1$.
@@
Total $4NM^{2}C^{3}+2NM^{4}C^{2}$ $(1+2\tau+3\tau^{2})NM^{2}C^{3}+2\tau^{2}NM^{4}C^{2}$ $(2\tau+3\tau^{2})NM^{2}C^{3}+2\tau^{2}NM^{4}C^{2}$
III-D Frequency Domain Motivation of the DCT-based Initialization